OpenStudy (anonymous):

a=-2 , b=-1 , c=4 , d=3 Solve) (Square Root Of:) -2b-3b+4=?

OpenStudy (anonymous):

are you solving for b?

OpenStudy (anonymous):

its like this : √-2(-1)-3(-1)+4=?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

you would have -2+3+4v in the end

OpenStudy (callisto):

@wickey1994 Disagree...

OpenStudy (anonymous):

om sorry i screwed that up, there should not be a negative in front of the 2 either,

OpenStudy (anonymous):

idk why is 3

OpenStudy (anonymous):

we didnt use a ,c and d

OpenStudy (callisto):

\[\sqrt{-2b-3b+4} = \sqrt{-2(-1)-3(-1)+4} =\sqrt{2+3+4} = \sqrt{9} = \pm3\]